# Matching Free Depth Recovery from Structured Light

### v1

@@ 5 Conclusion
In this paper, we proposed a novel framework for depth reconstruction in structured light systems using 3D voxel grids. Our approach centered on training a density voxel grid to represent the geometry of the captured scene, leveraging constraints from projected patterns to guide the training schedule through a fully differentiable rendering process. Upon convergence, we used volume density queried from the trained voxel grid to obtain a depth map through similar rendering approach. The key advantage of our method is that it eliminates the need for traditional correspondence searching on images, thereby removing the reliance on the accuracy of image matching algorithms. Experimental results show that our method yields competitive results compared to traditional matching-based algorithms, while using same or less amount of patterns as theirs. When compared with similar rendering-based method, which use implicit function to represent geometries, our approach demonstrates superior depth estimation accuracy while having faster training speed.
Acknowledgement This research was supported by the National Natural Science Foundation of China (No.62441224, No.62272433), the Fundamental Research Funds for the Central Universities (No.WK0010000090). The numerical calculations in this paper have been done on the supercomputing system in the Supercomputing Center of University of Science and Technology of China.

### v2

@@ V Conclusion
In this paper, we propose a novel framework for depth recovery in structured light systems using 3D voxel grids. Our approach centered on training a density voxel grid to represent the geometry of the captured scene, leveraging constraints from projected patterns to guide the training schedule through a fully differentiable rendering process. Upon convergence, we used volume density queried from the trained voxel grid to obtain a depth map through a similar rendering approach. A key advantage of our approach is that it completely eliminates the need for traditional correspondence search in the image space, thereby avoiding dependence on potentially error-prone matching algorithms. Experimental results demonstrate that our method achieves competitive, and in many cases superior, performance compared to conventional matching-based approaches and deep-learning-based techniques, while requiring the same or even fewer projected patterns. When compared with similar rendering-based methods, which use implicit functions to represent geometries, our approach demonstrates superior depth estimation accuracy while having faster training speed.